### v3

@@ 4 A structural model for ownership chains
Motivated by previous empirical evidence, we present a simple model to explain location strategies along ownership chains in multinational enterprises (MNEs). Inspired by Head and Ries (2008), we assume a competitive contest in which headquarters compete to invest in new subsidiary assets worldwide while incurring monitoring costs. Specifically, in the case of ownership chains, each headquarters faces a choice: it can either directly monitor a final target subsidiary or delegate monitoring to an intermediate managerial unit - an intermediate affiliate (sometimes referred to as a middleman). These intermediate affiliates are subsidiaries that are already part of the corporate structure before the auction takes place. A subsidiary is added to an ownership chain when the headquarters delegates or self-delegates monitoring, i.e. when the cost of delegation is effectively zero. In the latter case, we have the special case of direct control.
@@
The investment decision process unfolds in two steps. First, the headquarters bids in an auction for a new target subsidiary, conditional on the inspection costs given the location of the target. Second, if it wins at the auction, it delegates the unit located in an intermediate jurisdiction that minimizes inspection costs. As a result, inspection costs affect the bidding strategy in the auction. Importantly, the headquarters always retains the option to monitor subsidiaries directly, in which case no intermediate affiliate is necessary.
@@
Our model abstracts from production processes between headquarters and subsidiaries. While our setup does not exclude the possibility of supply chain coordination, it also applies to conglomerates and diversified business portfolios, where monitoring remains crucial for governance and managerial control.161616 To assess whether the length of ownership chains is associated with increasing distance in the production space, we ran a regression in which the hierarchical distance of an affiliate from its parent served as the dependent variable, including, among other controls, the “production distance”, i.e. $upstreamness_{ij}$, a measure from Alfaro et al. (2019). While the coefficient was statistically significant, its magnitude was negligible. This suggests that an implausibly long ownership chain would be required for an economically meaningful difference in production stages between parent and subsidiary. In our data, these results suggest that ownership chains are primarily shaped by organizational considerations rather than sequential production optimization. Results are available upon request.
4.1 The inspection game
@@
To formalize the decision-making process, we introduce a classical inspection game between the headquarters (Hq) and a target subsidiary (Sub). This framework helps headquarters determine the maximum bid they can offer in auctions for new subsidiaries. The possible strategies and corresponding payoffs for both players are summarized in Table 5.171717A textbook version of this inspection game was originally introduced by Fudenberg and Tirole (1991) and later adapted to the FDI context by Head and Ries (2008). We adapt their framework by incorporating the possibility of an intermediate monitoring unit within the multinational firm’s ownership network.
@@
In this corporate governance game, profits arise from the value generated by the headquarters (Hq) and its existing perimeter of subsidiaries ($A$), [^18] The existence of a corporate perimeter before new acquisitions implies that $A=\sum_{f}a_{f}$, meaning that the total value is generated by both Hq and all existing subsidiaries, regardless of new acquisitions. as well as by the additional subsidiary ($b$), conditional on the effort exerted by the subsidiary’s manager ($e$). Headquarters can choose to monitor $q$ or trust $1-q$ the subsidiary. Similarly, the subsidiary’s manager can work ($1-p$) or shirk ($p$). If the subsidiary works, headquarters pays a wage $w$; however, if headquarters monitors and discovers shirking, the wage and the value of the subsidiary default to zero ($w=b=0$).
@@
$Work$ $Shirk$
$(1-p)$ $p$
Hq $Trust$ (1-q) $A+b-w,w-e$ $A-w,w$
$Monitor$ ($q$) $A+b-w-C,w-e$ $A-C,0$
The headquarters (Hq) bears an inspection cost ($C$) to verify the activities of the subsidiary (Sub), a concept that we elaborate on in the next section. In the classical version of this game, we assume $b\geq w\geq e\geq C\geq 0$, which results in no pure-strategy Nash equilibrium. Therefore, in a mixed-strategy equilibrium, the expected payoffs for Hq and Sub are:
$v=A+b(1-p)-Cq-w(1-pq)$
$u=w(1-pq)-e(1-p)$
The first-order conditions for Hq and Sub are given by:
$v_{q}=-C+wp=0\quad\text{and}\quad v_{p}=-wq+e=0$
From these, we derive the equilibrium mixing probabilities:
$p=C/W\quad\text{and}\quad q=e/w$
Substituting these into Hq’s expected payoff, we obtain an initial expression for the valuation of Sub by Hq:
$v=A+b(1-C/w)-w$
Substituting this into Equation 4.1, we obtain the final valuation function that Hq assigns to Sub:
@@
Since corporate control operates in a global market, all Hqs play the same inspection game when evaluating subsidiaries that are available for acquisition. Consequently, competition ensures that the winning Hq is the one offering the highest valuation, determined by equation 2. It is important to note that, compared to the classical literature on delegation and monitoring in hierarchical structures (Holmstrom and Tirole, 1989), our framework relies on a number of simplifying assumptions regarding the contractual environment. First, we assume no information asymmetry between Hqs and intermediate affiliates (Mids), i.e. each party perfectly observes the other’s effort. Second, we exclude the possibility of side contracting, thus eliminating potential collusion between Mids and Subs. Finally, we treat the effort made by Mids as exogenous. In the next section, we examine how these assumptions shape Hq’s decision process and monitoring cost design.
4.2 Monitoring and delegation costs
When acquiring a subsidiary (Sub), headquarters (Hq) can choose between two monitoring strategies: (i) direct monitoring, where Hq monitors the subsidiary itself; or (ii) delegated monitoring, where Hq delegates the monitoring task to an intermediate affiliate (Mid). There are costs associated with both monitoring and delegation. More specifically, a Hq in country $i$ that wants to monitor a Sub in country $j$ wants to minimize the following composite cost function:
@@ 4.2 Monitoring and delegation costs
where: $\delta_{kj}$ represents the cost incurred by a Mid in country $k$ to monitor a Sub in country $j$; $\delta_{ik}$ represents the cost incurred by Hq in country $i$ to delegate monitoring to a Mid in country $k$; $\epsilon_{ikj}$ is a stochastic unobservable factor specific to the $ikj$-th ownership chain that captures idiosyncratic variations in the business environment that may increase or decrease monitoring costs. Notably, we assume, realistically, that delegated monitoring is always less expensive than direct monitoring, such that: $\delta_{kj}\geq\delta_{ik}$>00.
@@
4.2.1 The case of domestic subsidiaries
Domestic subsidiaries require a separate discussion because they introduce three different location strategies within ownership chains:
@@ 4.2.1 The case of domestic subsidiaries
    All three entities — Hq, Mid, and Sub — are located within the same home country ($i$). In this case, the composite cost function simplifies to: $02\delta_{ii}>$ 0. Here, monitoring and delegation costs are identical but remain greater than zero because domestic frictions – although smaller than foreign frictions – are still present. Since internal distances differ between countries, geography alone does not determine the chain structure. In the absence of other cost factors, Mid and Sub could theoretically swap positions within the ownership chain without affecting total costs, meaning that only the sequence of acquisitions determines their relative placement.
@@
4.2.2 The case of direct control
The simplest strategy for Hq is to retain full corporate control and not delegate monitoring to a Mid. In this case, Hq self-delegates, which means that delegation costs are effectively zero. More formally, in our model, direct control is treated as a valid location choice without delegation, represented as a $ikj$-th triplet where the intermediate jurisdiction ($k$) is null, i.e. $k=\emptyset$.
@@ 4.3.1 Locations of the chain
$=\mathds{P}(\delta_{ik}+\delta_{kj}-\epsilon_{ikj}\leq\delta_{i\ell}+\delta_{\ell j}-\epsilon_{i\ell j},\forall\ell\neq k)$
$=\mathds{P}(\epsilon_{i\ell j}\leq\delta_{i\ell}+\delta_{\ell j}-\delta_{ik}-\delta_{kj}+\epsilon_{ikj},\forall\ell\neq k)$
which we can write:
[…]
where $\mathcal{M}_{ikj}=1$ if Hq in country $i$ willing to invest in country $j$ chooses its monitoring unit in country $k$. F[•] is the joint cumulative distribution function, conditional on the value of $\epsilon_{ikj}$. By integrating $F[\bullet]$ with respect to the marginal probability density function $f(\epsilon_{ikj})$, we obtain the unconditional probability - the probability that country $k$ is chosen as the monitoring location for any given realization of $\epsilon_{ikj}$.
Since $f(\epsilon_{ikj})$’s are iid, the conditional probability is the product of $K-1$ univariate cumulative density functions:
$\mathds{P}(\mathcal{M}_{ikj}=1)$ $=\int_{-\infty}^{+\infty}\prod_{\ell\neq k}F_{\ell}[\delta_{i\ell}+\delta_{\ell j}-\delta_{ik}-\delta_{kj}+\epsilon_{ikj}]f(\epsilon_{ikj})\,d\epsilon_{ikj}$ (5)
$=\int_{-\infty}^{+\infty}e^{-\epsilon_{ikj}}exp(-e^{-\epsilon_{ikj}})\prod_{\ell\neq k}exp(-exp(-(\delta_{i\ell}+\delta_{\ell j}-\delta_{ik}-\delta_{kj}+\epsilon_{ikj}))\,d\epsilon_{ikj}$
Using the change of variable $z_{k}=e^{-\epsilon_{ikj}}$ and $y_{\ell}=e^{-(\delta_{i\ell}+\delta_{\ell j})}$ integrating by substitution:
$\mathds{P}(\mathcal{M}_{ikj}=1)$ $=\int_{0}^{+\infty}e^{-z_{k}}\prod_{\ell\neq k}e^{-\frac{y_{\ell}}{y_{k}}z_{k}}\,dz_{k}$
[…]
$=-\frac{y_{k}}{\sum\limits_{\ell}y_{\ell}}\biggr{[}e^{-z_{k}\sum\limits_{\ell=1}^{N}\frac{y_{\ell}}{y_{k}}}\biggr{]}_{0}^{+\infty}$
$=\frac{y_{k}}{\sum\limits_{\ell}y_{\ell}}$
$=\frac{e^{-(\delta_{ik}+\delta_{kj})}}{\sum\limits_{\ell}e^{-(\delta_{i\ell}+\delta_{\ell j})}}$
@@ 4.3.2 Probability of winning the auction for corporate control
$=\mathds{P}(a_{n}^{max}\leq a_{i}^{max}-2\sqrt{bC_{ij}}+2\sqrt{bC_{nj}},\forall n\neq i)$
where $\mathcal{N}_{ij}=1$ if an Hq in country $i$ wins control of a Sub in country $j$. We denote by $m_{i}$ the number of Hqs in country $i$. $A_{i}$ is distributed as a Gumbel with parameters $\mu$ and $\sigma$ and the maximum of $m$ Gumbel draws is distributed as a Gumbel with the same $\sigma$ and $\mu$ right-shifted by a quantity $\sigma\ln(m)$.
Following the same steps as in equations 4 and 5, and integrating by substitution, we derive the second testable equation.
@@ 5 Empirical strategy
In this section, we connect theory and empirics to identify the structural parameters of equations 6 and 7. The first equation describes the parent’s strategy for locating new subsidiaries along ownership chains, while the second characterizes the probability that the parent will gain control of a subsidiary in a given location.
We assume that $\pi_{ik\mid j}$ is identical across parents, which allows us to derive an aggregate form of the testable equation. By summing individual location choices, we define $M_{ikj}$, which represents the number of parents in country $i$ that delegate monitoring responsibilities to an intermediate affiliate in country $k$ for final subsidiaries located in country $j$ [^20] Note that direct control is a special case of an ownership chain in which the parent self-delegates monitoring. In this case, $k$ represents a null country and the delegation cost is zero. We obtain the expected value of $M_{ikj}$ as:
@@
Given the previous empirical equation, we now provide an interpretation for the costs of delegation ($\delta_{ik}$) and monitoring ($\delta_{kj}$), building on the evidence presented in Section 3 and Appendix B. Following the arguments of Stein and Daude (2007) and Bahar (2020), we proxy for organizational frictions using communication barriers between the country of the parent ($i$) and the country of the intermediate affiliate ($k$), and between the country of the intermediate affiliate ($k$) and the country of the ultimate subsidiary ($j$). We assume that time zone differences serve as an important source of friction, as they make it difficult to communicate in real time along the chain of ownership. Considering a 10-hour working day, we compute the number of overlapping working hours between firms based on their respective country locations. Thus, we denote by $wh_{ik}$ and $wh_{kj}$ the number of overlapping hours between the parent and the intermediate affiliate on the one hand, and between the intermediate affiliate and the final subsidiary on the other hand. Using this framework, we derive the following structural triangular equation:
$\mathds{E}\biggr{[}\frac{M_{ikj}}{M_{ij}}\biggr{]}=exp(\beta wh_{ik}+\rho wh_{kj}+\alpha_{1}X_{ik}+\alpha_{2}X_{kj}+FE_{ij})$ (8)
where the result represents the share of triplets ($ijk$) among all observed triplets, and $\beta$ and $\rho$ represent our coefficients of interest. Beyond organizational frictions, we introduce two vectors of geographic and institutional control variables - $X_{ik}$ and $X_{kj}$ that account for additional factors that may hinder coordination along ownership chains. These controls include bilateral geographic distance, common borders, common language, common legal origin, colonial ties, and participation in regional trade agreements. Finally, we include the bilateral fixed effect ($FE_{ij}$), which captures the expected aggregate cost index, $\mathcal{C}_{ij}$, as from the denominator of Equation 6.
@@
where $\theta$ is our coefficient of interest, and $FE_{i}$ and $FE_{j}$ represent origin and destination fixed effects, respectively. Finally, equations 8 and 9 form a system of two testable equations, which we estimate using a Pseudo-Poisson Maximum Likelihood (PPML) estimator, following Silva and Tenreyro (2006).
@@ 6 Results
The baseline results are reported in Table 6. The first column presents estimates for the triangular equation specified in equation 8. Our results confirm that ease of communication has a significant impact on delegation and monitoring decisions, thereby influencing the location choices of intermediate affiliates. An increase of one overlapping working hour between the parent and intermediate affiliates increases the expected fraction of final subsidiaries controlled by the intermediate $k$ by 5.6%, calculated as $(e^{0.055}-1)\times 100$. Similarly, an increase of one overlapping work hour between the intermediary and the location of the final subsidiary increases the expected share of subsidiaries controlled by the intermediate $k$ by 10.5%, calculated as $(e^{0.1}-1)\times 100$. Importantly, the relative magnitudes of these effects match expectations. As anticipated, delegation costs are less binding than monitoring costs, meaning that a reduction in the second coefficient (monitoring) has about twice the effect of a reduction in the first (delegation).
Table 6: Baseline results
Location of: Middlemen Final subsidiaries
Dep. var. $M_{ikj}/M_{ij}$ $M_{ij}$
@@
[…]
$\text{RTA}_{ik}$ 1.261***
$\text{RTA}_{kj}$ 0.807***
$\text{Language}_{ik}$ 0.685***
$\text{Language}_{kj}$ 0.419***
$\text{Colony dependence}_{ik}$ 1.534***
$\text{Colony dependence}_{kj}$ 1.520***
$\text{Legal origins}_{ik}$ 0.236***
$\text{Legal origins}_{kj}$ 0.225***
$\text{Home}_{ik}$ 5.392***
$\text{Home}_{kj}$ 3.597***
$\text{Direct control}_{ikj}$ 0.836***
$\sqrt{\frac{\hat{\mathcal{C}}_{ij}}{K}}$ –140.892***
Observations 1,718,840 9,712
Fixed effects $i\times j$ i,j
  • Note: Given the presence of zeros in the case of direct control, we scale the distances in km by using the hyperbolic sine transformation ($ln(x+\sqrt{x^{2}+1})$), which allows approximating to the natural logarithm while retaining the zeros (Bellemare and Wichman, 2020). Standard errors are double-clustered by origin and destination in parentheses (*** p$<$0.01, ** p$<$0.05, * p$<$0.1).
In addition to organizational costs, we assumed that other controls capture both geographic and institutional frictions that affect delegation and monitoring decisions. Interestingly, the coefficients on geographic distance follow a pattern similar to that of overlapping hours. An increase in the distance between the parent and the intermediate affiliate reduces the share of subsidiaries controlled by country $k$ by 2.9% ($(e^{0.03}-1)\times 100$). Similarly, increasing the distance between the intermediate firm and the subsidiary reduces the share of subsidiaries controlled by country $k$ by 15.8% ($(e^{0.172}-1)\times 100$).
Finally, the last column of Table 6 presents the coefficient of the vector of the expected cost of corporate control ($E[C_{ikj}]$), obtained from previous estimates. The coefficient obtained corresponds to $\theta$ in equation 9, and its sign and magnitude are as expected: an increase in the cost index reduces the number of subsidiaries in country $j$ controlled by parents in country $i$. Figure 8 shows the distribution of the expected inspection cost for each origin $i$ and destination $j$, which is a useful byproduct of our structural model.

### v4

@@ 4 An analytical framework for ownership chains
In this section, we provide a conceptual framework to explain location strategies along ownership chains in multinational enterprises (MNEs). We extend the Head and Ries (2008) model to incorporate the possibility of ownership chains. The original Head and Ries (2008)’ setup assumes that headquarters compete to win control over new subsidiary assets worldwide through an auction, while incurring monitoring costs. In the context of ownership chains, we posit that each headquarters faces a choice: it can either directly monitor a final target subsidiary or delegate monitoring to an intermediate managerial unit, an intermediate affiliate (sometimes referred to as a middleman). These intermediate affiliates are subsidiaries that are already part of the corporate structure prior to the auction taking place  [^12] In principle, intermediate affiliates could be monitored themselves. However, incorporating recursive monitoring would complicate the model without generating additional testable implications for the initial allocation of control along the ownership chain. Therefore, within the present framework, intermediate affiliates are best understood as organizational nodes to which monitoring authority is delegated, rather than as agents whose incentives are the primary focus of this study. More generally, delegation of monitoring authority is itself a meaningful organizational margin, even when monitoring technologies are identical, because it reshapes incentive and commitment constraints (Strausz, 1997).. A subsidiary is added to an ownership chain when the headquarters delegates or self-delegates monitoring, i.e., when the cost of delegation is effectively zero. In the latter case, we have the special case of direct control.
@@
We conceptualize the investment decision process as unfolding in two stages. First, the headquarters bids in an auction for a new target subsidiary, contingent upon the inspection costs given the target’s location. Second, if it wins the auction, it delegates the unit located in an intermediate jurisdiction that minimizes inspection costs. As a result, inspection costs affect the bidding strategy in the auction. Importantly, headquarters always retains the option to monitor subsidiaries directly, in which case no intermediate affiliate is necessary.
@@
Note that the model underlying our analysis abstracts from production processes between headquarters and subsidiaries. While our setup does not exclude the possibility of supply chain coordination, it also applies to conglomerates and diversified business portfolios, where monitoring remains crucial for governance and managerial control. [^13] To assess whether the length of ownership chains is associated with increasing distance in the production space, we ran a regression in which the hierarchical distance of an affiliate from its parent served as the dependent variable, including, among other controls, the “production distance”, i.e. $upstreamness_{ij}$, a measure from Alfaro et al. (2019). While the coefficient was statistically significant, its magnitude was negligible. This suggests that an implausibly long ownership chain would be required for an economically meaningful difference in production stages between parent and subsidiary. In our data, these results suggest that ownership chains are primarily shaped by organizational considerations rather than sequential production optimization. Results are available upon request.
@@
Head and Ries (2008) formalize the decision-making process introducing a classical inspection game between the headquarters (Hq) and a target subsidiary (Sub). This framework helps headquarters determine the maximum bid they can offer in auctions for new subsidiaries. The possible strategies and corresponding payoffs for both players are summarized in Table 5.141414A textbook version of this inspection game was originally introduced by Fudenberg and Tirole (1991) and later adapted to the FDI context by Head and Ries (2008). We adapt their framework by incorporating the possibility of an intermediate monitoring unit within the multinational firm’s ownership network.
@@
In this corporate governance game, profits arise from the value generated by the headquarters (Hq) and its existing perimeter of subsidiaries (AA) [^15] The existence of a corporate perimeter before new acquisitions implies that $A=\sum_{f}a_{f}$, meaning that the total value is generated by both Hq and all existing subsidiaries, regardless of new acquisitions., as well as by the additional subsidiary ($b$), conditional on the effort exerted by the subsidiary’s manager ($e$). Headquarters can choose to monitor the subsidiary with probability $q$ or trust it with probability 1q1-q . Similarly, the subsidiary’s manager can work (1p1-p) or shirk ($p$). If the subsidiary works, headquarters pays a wage $w$; however, if headquarters monitors and discovers shirking, the wage and the value of the subsidiary default to zero ($w=b=0$).
@@
WorkWork ShirkShirk
(1p)(1-p) $p$
Hq TrustTrust (1-q) A+bw,weA+b-w,w-e Aw,wA-w,w
MonitorMonitor ($q$) A+bwC,weA+b-w-C,w-e AC,0A-C,0
The headquarters (Hq) bears an inspection cost ($C$) to verify the activities of the subsidiary (Sub), a concept that we elaborate on in the next section. In the classical version of this game, we assume $b\geq w\geq e\geq C\geq 0$, which results in no pure-strategy Nash equilibrium. Therefore, in a mixed-strategy equilibrium, the authors obtain a final valuation function that Hq assigns to Sub:
@@
Since corporate control operates in a global market, all Hqs play the same inspection game when evaluating subsidiaries that are available for acquisition. Consequently, competition ensures that the winning Hq is the one offering the highest valuation, determined by equation 3.
With respect to the original model, we assume that, when acquiring a subsidiary (Sub), headquarters (Hq) can choose between two monitoring strategies: (i) direct monitoring, where Hq monitors the subsidiary itself; or (ii) delegated monitoring, where Hq delegates the monitoring task to an intermediate affiliate (Mid). There are costs associated with both monitoring and delegation. More specifically, a Hq in country $i$ that wants to monitor a Sub in country $j$ wants to minimize the following composite cost function:
@@
where: $\delta_{kj}$ represents the cost incurred by a Mid in country $k$ to monitor a Sub in country $j$; $\delta_{ik}$ represents the cost incurred by Hq in country $i$ to delegate monitoring to a Mid in country $k$; $\epsilon_{ikj}$ is a stochastic unobservable factor specific to the ikjikj-th ownership chain that captures idiosyncratic variations in the business environment that may increase or decrease monitoring costs. Notably, we assume, realistically, that delegated monitoring is always less expensive than direct monitoring, such that: $\delta_{kj}\geq\delta_{ik}$>0.
@@
Domestic subsidiaries require a separate discussion because they introduce three different location strategies:
@@
    All three entities — Hq, Mid, and Sub — are located within the same home country ($i$). In this case, the composite cost function simplifies to: $2\delta_{ii}>0$. Here, monitoring and delegation costs are identical but remain greater than zero because domestic frictions – although smaller than foreign frictions – are still present. Since internal distances differ between countries, geography alone does not determine the chain structure. In the absence of other cost factors, Mid and Sub could theoretically swap positions within the ownership chain without affecting total costs, meaning that only the sequence of acquisitions determines their relative placement.
@@
Finally, our framework also accounts for directly controlled subsidiaries. In fact, the simplest strategy for Hq is to retain full corporate control and not delegate monitoring to a Mid. In this case, Hq self-delegates, which means that delegation costs are effectively zero. More formally, in our model, direct control is treated as a valid location choice without delegation, represented as a ikjikj-th triplet where the intermediate jurisdiction ($k$) is null, i.e., $k=\emptyset$.
@@ 4.1.1 Locations of the chain
After some algebraic derivations, we obtain the following result [^17] All intermediate steps are available upon request.
$\mathds{P}(\mathcal{M}_{ikj}=1)=\frac{e^{-(\delta_{ik}+\delta_{kj})}}{\sum\limits_{\ell}e^{-(\delta_{i\ell}+\delta_{\ell j})}}\\$
@@ 4.1.2 Probability of winning the auction for corporate control
$\mathds{P}(a_{n}^{max}\leq a_{i}^{max}-2\sqrt{bC_{ij}}+2\sqrt{bC_{nj}},\forall n\neq i)$
where $\mathcal{N}_{ij}=1$ if an Hq in country $i$ wins control of a Sub in country $j$. We denote by $m_{i}$ the number of Hqs in country $i$. After some algebraic derivations, we derive the second testable equation [^18] All intermediate steps are available upon request..
@@ 5 Empirical strategy
Equations 5 and 6 constitute the foundation of our empirical strategy. The first equation describes the parent’s strategy for locating new subsidiaries along ownership chains, while the second characterizes the probability that the parent will gain control of a subsidiary in a given location.
We assume that $\pi_{ik\mid j}$ is identical across parents, which allows us to derive an aggregate form of the testable equation. By summing individual location choices, we define $M_{ikj}$, which represents the number of parents in country $i$ that delegate monitoring responsibilities to an intermediate affiliate in country $k$ for final subsidiaries located in country jj. [^19] Note that direct control is a special case of an ownership chain in which the parent self-delegates monitoring. In this case, $k$ represents a null country and the delegation cost is zero. We obtain the expected value of $M_{ikj}$ as:
@@
From the above empirical equation, we derive this triangular gravity equation:
$\mathds{E}\biggr[\frac{M_{ikj}}{M_{ij}}\biggr]=exp(\alpha_{1}X_{ik}+\alpha_{2}X_{kj}+FE_{ij})$ (7)
where the result represents the share of triplets (ijkijk) among all observed triplets, and $\beta$ and $\rho$ represent our coefficients of interest. $X_{ik}$ and $X_{kj}$ are vectors of geographic and institutional control variables that capture factors potentially hindering coordination along ownership chains.
Beyond standard variables, such as bilateral distance, contiguity, common language, shared legal origin, colonial ties, and participation in regional trade agreements, these vectors also embed a key measure of organizational frictions: the number of overlapping working hours.
Building on Stein and Daude (2007) and Bahar (2020), we interpret time-zone misalignment as a salient barrier to real-time communication. Limited overlap in working hours between the parent country $i$ and the intermediate affiliate $k$, as well as between $k$ and the final subsidiary $j$, raises coordination costs and can significantly shape the structure of multinational ownership chains. Finally, we include the bilateral fixed effect ($FE_{ij}$), which captures the expected aggregate cost index, $\mathcal{C}_{ij}$, as from the denominator of Equation 5.
@@
where $\theta$ is our coefficient of interest, and $FE_{i}$ and $FE_{j}$ represent origin and destination fixed effects, respectively. Finally, equations 7 and 8 form a system of two testable equations, which we estimate using a Pseudo-Poisson Maximum Likelihood (PPML) estimator, following Silva and Tenreyro (2006).
@@ 6 Results
Table 6 reports estimates for equations 7 and 8 across three nested specifications, moving from conventional gravity predictors to variables more closely related to cross-border communication and coordination.
Columns (1), (3) and (5) show results for the triangular equation specified in equation 7. In the first column we present a restricted specification that includes geographical distance (measured in kilometers), a dummy for the presence of a regional trade agreement (RTA), and a dummy indicating domestic control relationships. We see that the farther the middleman is from either the parent or the final destination, the smaller the share of investments routed through $k$, with the effect being stronger for distance to the final destination ($26\%$) than for distance to the parent ($5.6\%$). Both the RTA and the domestic-control dummies have positive and statistically significant effects.
Column (3) augments the model with standard gravity variables that capture information, legal, and regulatory costs, including common language, common legal origin, and colonial ties. These variables are positively associated with the share of subsidiaries controlled by an intermediary in country k and, with the exception of common language, have similar effects in both delegation relationships and monitoring relationships. When the parent and middleman countries speak the same language, the share variable increases by 98%, when it is the final country to speak the same language as the middleman, the increase is 55%. Coefficients on shared predictors remain significant and quite stable between Columns (1) and (3).
Column (5) presents the full specification, which introduces the number of overlapping working hours between countries as a direct measure of frictions to real-time communication. This predictor has a significant impact on delegation and monitoring decisions. An increase of one overlapping working hour between the parent and intermediate affiliates increases the expected fraction of final subsidiaries controlled by the intermediate $k$ by 5.6%. Similarly, an increase of one overlapping work hour between the intermediary and the location of the final subsidiary increases the expected share of subsidiaries controlled by the intermediate $k$ by 10.5%.
Importantly, the relative magnitudes of these effects match expectations. As anticipated, delegation costs are less binding than monitoring costs, meaning that a reduction in the second coefficient (monitoring) has about twice the effect of a reduction in the first (delegation). This is in line with our assumptions about Equation 4.
The inclusion of overlapping working hours also attenuates the estimated effects of geographical distance. Relative to Column (3), the distance coefficients decrease by nearly half, suggesting a positive correlation between distance and overlapping working hours. This pattern suggests that, in earlier specifications, geographical distance partially captures coordination frictions that are more directly measured by the overlapping-hours introduced in Column (5).
Table 6: Structural gravity results
(1) (2) (3) (4) (5) (6)
Location of: Middlemen Final subsidiaries Middlemen Final subsidiaries Middlemen Final subsidiaries
Dep. var. $M_{ikj}/M_{ij}$ $M_{ij}$ $M_{ikj}/M_{ij}$ $M_{ij}$ $M_{ikj}/M_{ij}$ $M_{ij}$
@@
$\text{Language}_{ik}$ 0.684*** 0.685***
(0.0318) (0.0315)
$\text{Language}_{kj}$ 0.441*** 0.419***
(0.0326) (0.0327)
$\text{Colony dependence}_{ik}$ 1.513*** 1.534***
(0.0483) (0.0479)
$\text{Colony dependence}_{kj}$ 1.503*** 1.520***
(0.0481) (0.0479)
$\text{Legal origins}_{ik}$ 0.231*** 0.236***
(0.0280) (0.0280)
$\text{Legal origins}_{kj}$ 0.221*** 0.225***
(0.0261) (0.0262)
$\text{Distance (arcsin)}_{ik}$ (km) –0.0575*** –0.0615*** –0.0302***
(0.00908) (0.00922) (0.00988)
$\text{Distance (arcsin)}_{kj}$ (km) –0.306*** –0.299*** –0.172***
(0.0101) (0.0101) (0.0125)
$\text{RTA}_{ik}$ 1.408*** 1.289*** 1.261***
(0.0271) (0.0275) (0.0276)
$\text{RTA}_{kj}$ 0.930*** 0.860*** 0.807***
(0.0251) (0.0257) (0.0258)
$\text{Home}_{ik}$ 5.319*** 5.432*** 5.392***
(0.0471) (0.0494) (0.0484)
$\text{Home}_{kj}$ 3.303*** 3.438*** 3.597***
(0.0596) (0.0592) (0.0571)
$\text{Direct control}_{ikj}$ 0.696*** 0.665*** 0.836***
(0.0518) (0.0528) (0.0565)
$\sqrt{\frac{\hat{\mathcal{C}}_{ij}}{K}}$ –156.576*** –154.541*** –140.892***
(16.517) (12.986) (14.004)
Observations 1,724,334 9,736 1,718,840 9,712 1,718,840 9,712
Fixed effects $i\times j$ i,j $i\times j$ i,j $i\times j$ i,j
  • Note: Given the presence of zeros in the case of direct control, we scale the distances in km by using the hyperbolic sine transformation ($ln(x+\sqrt{x^{2}+1})$), which allows approximating to the natural logarithm while retaining the zeros (Bellemare and Wichman, 2020). Standard errors are double-clustered by origin and destination in parentheses (*** p<<0.01, ** p<<0.05, * p<<0.1).
Columns (2), (4) and (6) of Table 6 present the coefficients of the vector of the expected cost of corporate control ($E[C_{ikj}]$), obtained from previous estimates. The coefficients obtained correspond to $\theta$ in equation 8, and their sign and magnitude are as expected: an increase in the cost index reduces the number of subsidiaries in country $j$ controlled by parents in country $i$. Finally, Figure 6 shows the distribution of the expected inspection cost for each origin $i$ and destination $j$ obtained from the full model in column (5).